# Q-score Max-Clique: The First Quantum Metric Evaluation on Multiple Computational Paradigms

### v2

@@ Abstract
Evaluating the performance of quantum devices is an important step towards scaling quantum devices and eventually using them in practice. The great number of available quantum metrics and the different hardware technologies used to develop quantum computers complicate this evaluation. In addition, different computational paradigms implement quantum operations in different ways. We add to the landscape of quantum metrics by extending the Q-score metric of Atos to the Q-score Max-Clique. To our knowledge, this yields the first application-level metric which allows comparison of three different paradigms of quantum computing. This metric is evaluated on these computational quantum paradigms – quantum annealing, gate-based quantum computing, and photonic quantum computing – and the results are compared to those obtained by classical solvers.
1 Introduction
@@
While comparing quantum devices based on the same technology is already hard, comparing quantum devices based on different technologies is a research field on its own. Quantum metrics provide a first step towards quantifying the performance of quantum devices in a unified manner. However, just as with quantum computers, a single best metric does not exist. Instead, there exist multiple metrics and each has its own benefits and limitations.
@@
Application-level metrics address these issues by measuring the performance of quantum devices on quantum algorithms and practical problems. An example is given by the suite of the QED Consortium [10], which considers a few standard quantum algorithms to compare the performance of quantum devices. While this application-level metric gives insight in the performance of the quantum device, it is still only applicable to gate-based devices. As gate-based quantum computers have the promise to be universal [11, 12], this is not unreasonable. However, other quantum computational paradigms such as photonic quantum computing and quantum annealing are also relevant and useful for specific problems.
This issue is solved by the Q-score metric, which considers the real-world Max-Cut problem to benchmark devices. The main idea behind the Q-score metric is that a quantum routine usually is part of a larger computational workflow, and benchmarks should therefore consider the whole computational pipeline instead of only the quantum part. For this reason, it is defined in a computing-agnostic way, because of which it can be used to benchmark any machine that can solve the Max-Cut problem [13]. Earlier work shows how the Q-score can be used to benchmark quantum annealers and classical solvers [14].
In this work, the original formulation of the Q-score is extended to define the Q-score Max-Clique, which allows comparison of quantum annealers, gate-based and photonic quantum devices. The metric is used to benchmark these three computational paradigms, as well as two classical solvers. To our knowledge, this is the first time performance of these four computational paradigms is compared.
The structure of this paper is as follows: In Section 2 background information is provided on quantum computing for the different technology paradigms and earlier Q-score work. Section 3 defines the Q-score Max-Clique metric and Section 4 presents the experiments and subsequent results. This work concludes with a discussion of the results and directions for further research in Section 5.
2 Background
2.1 Quantum Computing
@@
2.1.1 Quantum annealing
Quantum annealing is based on a different form of quantum computing called adiabatic quantum computing [15], which is known to be equivalent to standard gate-based quantum computing [16]. In adiabatic quantum computing qubits are not addressed individually, but, instead, qubits are prepared in some initial state and a certain Hamiltonian is slowly applied to these qubits. The Hamiltonian is constructed so that its ground-state encodes the solution of the problem. If this evolution is slow enough, the state remains in the ground state of the Hamiltonian throughout evolution, yielding the solution to the problem. The system is required to remain coherent during the evolution time. As this evolution time of the process can be exponential in the size of the problem, there are certain limitations to adiabatic quantum computing.
@@ 2.1.1 Quantum annealing
2.1.2 Gate-based quantum computing
@@
2.1.3 Photonic quantum computing
@@
2.2 Q-score
In 2021, the company Atos proposed the original Q-score metric with the idea that the state-of-the-art benchmarking of quantum devices was insufficient to assess High-Performance Computing platforms [13]. They designed the Q-score with the idea of having a metric which is: “application-centric, hardware-agnostic and scalable to quantum advantage processor sizes and beyond".
@@ 2.2 Q-score
From the definition of $\beta(N)$ it follows that a random approach results in a $\beta(N)$ of roughly zero. In contrast, an algorithm that always finds the best possible cut will give $\beta(N)\approx 1$. The value $\beta(N)$ hence captures how ‘good’ a cut is, with higher $\beta(N)$ values corresponding to better cuts. The Q-score of a device is defined as the largest $N$ for which $\beta(N)>\beta^{*}$, for some threshold value $\beta^{*}$. Atos suggest a threshold value $\beta^{*}=0.2$ to represent a successful execution of the algorithm. Thus, any algorithm achieving a $\beta(N)$ value higher than $0.2$ successfully solved the problem. Other choices might suffice as well, however $\beta^{*}=0.2$ is chosen so that a noiseless gate-based quantum device running the Quantum Approximate Optimisation Algorithm (QAOA) [22] with $p=1$, obtains an infinite Q-score.
@@
The original definition of the Q-score allows the user to choose some variables, such as the allowed time for the computation. This impacts the Q-score significantly, as classical devices can solve the problem for any $N$ perfectly, given a sufficient amount of time. Other examples of user variables to choose are the number of random instances of Erdös-Rényi graphs and the allowed pre- and post-processing optimisation steps. Earlier work [14] imposed a time-constraint of at most $60$ seconds for each instance, a total of $100$ random graph instances per graph size and considered each solver in an out-of-the-box fashion to restrict any pre- or post-processing optimisation steps. If a device failed to find an answer within $60$ seconds, a random cut was used for that instance, corresponding to a value of $\beta(N)=0$. This earlier work showed that, given these constraints, classical backends achieve higher Q-scores ($2{,}300$ and $5{,}800$) than state-of-the-art quantum annealing hardware ($70$ and $140$). In turn, the annealing-classical hybrid method offered by D-Wave Systems Inc. achieved a Q-score of $12{,}500$, which hints at gains offered by quantum annealing. The work of [14] also briefly discusses gate-based devices and their expected lower Q-score, mainly due to their low number of available qubits, and does not discuss photonic devices.
3 The Q-score Max-Clique metric
To extend upon the original Q-score metric, the Q-score Max-Clique is introduced: a new and improved version of the Q-score. This new metric takes the ideas of the Q-score and enhances them so that the metric can also be evaluated on photonic quantum devices, in addition to gate-based quantum devices, quantum annealers and classical devices.
@@ 3 The Q-score Max-Clique metric
Just like the Q-score, the Q-score Max-Clique aims to gauge the performance of a device by measuring the largest problem size $N$ of a certain optimisation problem for which the device significantly outperforms a random solution. The Max-Clique problem, contrary to the Max-Cut problem, allows for a more natural comparison between all mentioned backends. The Max-Clique problem is an established NP-hard problem in graph theory which aims to find the largest complete subgraph of a random graph of size $N$. A complete subgraph, or clique, is a subgraph in which every two vertices are connected by an edge. Just like the Max-Cut problem, the Max-Clique problem has a QUBO-formulation and the problem is hence equivalent to many optimisation problems, in particular to the Max-Cut problem [23]. The Q-score Max-Clique hence also serves as a metric on how well a device performs at general optimisation problems.
For computing the Q-score Max-Clique, the same variables as for the original Q-score are used: $(N,\frac{1}{2})$-Erdös-Rényi graphs are considered and the cost $C(N)$ is computed as the average clique size found over $100$ random instances of these graphs. The value $\beta(N)$ for this algorithm is then defined as
$\beta(N)=\frac{C(N)-C_{rand}}{C_{max}-C_{rand}}.$
@@
Here, $C_{rand}$ is the average clique size found by the following random naive approach: start with an empty sub-graph $G^{\prime}$; Take a random node $v\in G\setminus G^{\prime}$ and add this to $G^{\prime}$ if, $G^{\prime}\cup\{v\}$ is still a clique. If not, return the number of nodes in $G^{\prime}$ as the maximum clique size found by this random algorithm. For general Erdös-Rényi graphs $G(N,p)$, the probability that this naive algorithm finds a clique of size at least $i$ is
@@
For $p=\frac{1}{2}$, this approximately gives $C_{rand}=\mathbb{E}[X]=1.6416325$. Again, $C_{max}$ is the actual maximum clique size of the graph, averaged over the 100 random instances. According to [24], this value scales as
@@
Note that this is an asymptotic result for $N\to\infty$ and might not be a good approximation for smaller $N$.
A device significantly outperforms a random algorithm for some problem instance $N$ if $\beta(N)>\beta^{*}$, where again $\beta^{*}=0.2$ is taken, following [13]. So a device outperforms a random algorithm if
@@
The limit on the computation time is set to $60$ seconds per instance, as in [14]. If a device does not find an answer within the allocated 60 seconds, a clique size of $C_{rand}$ is assigned, corresponding to a value of $\beta(N)=0$. Finally, all algorithms are considered in an out-of-the-box fashion to restrict any additional optimisation steps.
4 Results
This section gives the Q-score Max-Clique results for different computational paradigms. First, a brief overview is given of how the Max-Clique problem can be solved on each of the quantum computing paradigms, as well as the devices for which the Q-score Max-Clique is computed.
@@ 4 Results
  • Gate-based quantum computing: The same QUBO formulation can be used to solve Max-Clique instances using the Quantum Approximate Optimisation Algorithm (QAOA). This is a hybrid quantum-classical algorithm in which a certain cost function, depending on measurements of quantum states, is optimised with a classical optimiser. For each optimisation step, a new quantum circuit needs to be run. Currently, quantum devices are accessed through the cloud, and resources are shared with other users. Because of this, each set of quantum circuits needs to be sent as a job through the entire stack and queue, which is rather inefficient, especially in variational settings, such as QAOA. Solutions are being developed that provide users the ability to avoid latency by executing entire programs in containerised services, such as Qiskit Runtime [25]. Still, significant queue times remain and once scheduled, a single problem instance can still take around an hour to execute in practice. The actual computation time QAOA would take with direct access to dedicated hardware is hard to estimate. A methodology to reason about the execution-time of QAOA is developed in [26]. According to their estimates, candidate solutions for Max-Cut graphs with $500$ nodes can be generated in under three minutes, which is still above the $60$ seconds threshold.
@@
In addition, the Q-score Max-Clique is also calculated for two classical solvers, namely qbsolv and simulated annealing, both offered by D-Wave Systems Inc. The qbsolv method solves smaller pieces of a QUBO with a tabu search algorithm [36, 37] and then combines those solutions to a full solution. The simulating annealing method is able to solve a QUBO by searching the solution space and slowly decreasing the probability of accepting worse solutions than the current optimum. Lastly, the Q-score Max-Clique is computed for the hybrid solver provided by D-Wave Systems Inc.. This is an algorithm which combines classical and annealing approaches to solve QUBOs. The precise workings of this solver are however unknown. Table 1 gives the Q-score of all considered backends.
Table 1: Q-scores Max-Clique with a $60$ seconds time limit. The starred entries were obtained using more than 60 seconds.
@@
Tabu search 4,900
Simulated annealing 9,100

### v3

# Extending the Q-score to an Application-level Quantum Metric Framework

@@ Abstract
Evaluating the performance of quantum devices is an important step towards scaling quantum devices and eventually using them in practice. The great number of available quantum metrics and the different hardware technologies used to develop quantum computers complicate this evaluation. In addition, different computational paradigms implement quantum operations in different ways. A prominent quantum metric is given by the Q-score metric of Atos. This metric was originally introduced as a standalone way to benchmark devices using the Max-Cut problem. In this work, we show that the Q-score defines a framework of quantum metrics, which allows benchmarking using different problems, user settings and solvers. To showcase the applicability of the framework, we showcase a second Q-score in this framework, called the Q-score Max-Clique. This yields, to our knowledge, the first application-level metric capable of natively comparing three different paradigms of quantum computing. This metric is evaluated on these computational quantum paradigms – quantum annealing, gate-based quantum computing, and photonic quantum computing – and the results are compared to those obtained by classical solvers.
Index Terms:
Quantum metrics, Q-score, Application-level benchmarking, Max-Clique
I Introduction
@@
While comparing quantum devices based on the same technology is already hard, comparing quantum devices based on different technologies is a research field on its own. Quantum metrics provide a first step towards quantifying the performance of quantum devices in a unified manner. However, just as with quantum computers, a single best metric does not exist. An overview of some of the leading metrics, as well as a classification of such metrics, is given by [6]. Instead, there exist multiple metrics and each has its own benefits and limitations.
@@
These issues are addressed by application-level metrics, which measure the performance of quantum devices on quantum algorithms and practical problems. There are various flavours of application-level metrics. Some consist of a suite of metrics considering different problems, such as the QED-C suite [11], SupermarQ [12] and QPack [13]. Others focus on a specific problem, such as the Q-score [14] and Quantum Linpack [15]. All of these metrics originally focus on gate-based devices. As gate-based quantum computers have the promise to be universal [16, 17], this is not unreasonable. However, other quantum computational paradigms such as photonic quantum computing and quantum annealing are also relevant and useful for specific problems.
This issue can be solved by extending the Q-score metric. The metric was originally defined as a stand-alone number considering the real-world Max-Cut problem to benchmark devices. The main idea behind the Q-score metric is that a quantum routine usually is part of a larger computational workflow, and benchmarks should therefore consider the whole computational pipeline instead of only the quantum part. For this reason, it can be defined in a computing-agnostic way, because of which it can be used to benchmark any machine that can solve the Max-Cut problem. Earlier work shows how the Q-score can be extended to benchmark quantum annealers and classical solvers [18] as well.
In this work, we show that the original formulation of the Q-score can be extended to form a quantum metrics framework. This is due to the fact that there are various parameters in the metric which can be freely chosen, most notably the optimisation problem. We showcase the strength of this degree of freedom by defining a second Q-score, called the Q-score Max-Clique. Our Max-Clique extension is specifically designed to yield a metric that can benchmark quantum annealers, gate-based and photonic quantum devices, as well as classical devices. This yields, to the best of our knowledge, the first metric capable of making a native comparison between these three paradigms. The metric is used to benchmark these three computational quantum paradigms, as well as two solvers based on classical computing. To our knowledge, this is the first time performance of these four computational paradigms is natively compared.
The structure of this paper is as follows: In Section II background information is provided on quantum computing for the different technology paradigms and earlier Q-score work. Section III explains how the Q-score can be adapted to define a framework, and details a second member of this framework called the Q-score Max-Clique metric. Afterwards, Section IV presents the experiments and subsequent results. This work concludes with a discussion of the results and directions for further research in Section V.
II Background
II-A Quantum Computing
@@
II-A1 Quantum annealing
Quantum annealing is based on adiabatic quantum computing [19], which is known to be equivalent to standard gate-based quantum computing [20]. In adiabatic quantum computing qubits are not addressed individually, but, instead, qubits are prepared in some initial state and a certain Hamiltonian is slowly applied to these qubits. The Hamiltonian is constructed so that its ground-state encodes the solution of the problem. If this evolution is slow enough, the state remains in the ground state of the Hamiltonian throughout evolution, yielding the solution to the problem. The system is required to remain coherent during the evolution time. As this evolution time of the process can be exponential in the size of the problem, there are certain limitations to adiabatic quantum computing.
@@ II-A1 Quantum annealing
II-A2 Gate-based quantum computing
@@
II-A3 Photonic quantum computing
@@
II-B Q-score
In 2021, the company Atos proposed the original Q-score metric with the idea that the state-of-the-art benchmarking of quantum devices was insufficient to assess High-Performance Computing platforms [14]. They designed the Q-score with the idea of having a metric which is: “application-centric, hardware-agnostic and scalable to quantum advantage processor sizes and beyond". Other strengths of the Q-score are that it is useful (it tells users what they should know), that it is comprehensive (it covers all relevant attributes) [24], and that it is easy to understand.
@@ II-B Q-score
From the definition of $\beta(N)$ it follows that a random approach results in a $\beta(N)$ of roughly zero. In contrast, an algorithm that always finds the best possible cut will give $\beta(N)\approx 1$. The value $\beta(N)$ hence captures how ‘good’ a cut is, with higher $\beta(N)$ values corresponding to better cuts. The Q-score of a device is defined as the largest $N$ for which $\beta(N)>\beta^{*}$, for some threshold value $\beta^{*}$. Atos suggest a threshold value $\beta^{*}=0.2$ to represent a successful execution of the algorithm. Thus, any algorithm achieving a $\beta(N)$ value higher than $0.2$ successfully solved the problem. Other choices might suffice as well, however $\beta^{*}=0.2$ was chosen so that a noiseless gate-based quantum device running the Quantum Approximate Optimisation Algorithm (QAOA) [27] with $p=1$, obtains an infinite Q-score.
@@
The original definition of the Q-score allows the user to choose some variables, such as the allowed time for the computation. This impacts the Q-score significantly, as classical devices would be able to solve the problem for any $N$ perfectly, resulting in an infinite Q-score, given a sufficient amount of time. Other examples of user variables to choose are the number of random instances of Erdös-Rényi graphs and the allowed pre- and post-processing optimisation steps. Earlier work [18] imposed a time-constraint of at most $60$ seconds for each instance, a total of $100$ random graph instances per graph size and considered each solver in an out-of-the-box fashion to restrict any pre- or post-processing optimisation steps. If a device failed to find an answer within $60$ seconds, a random cut was used for that instance, corresponding to a value of $\beta(N)=0$. This earlier work showed that, given these constraints, classical backends achieve higher Q-scores ($2{,}300$ and $5{,}800$) than state-of-the-art quantum annealing hardware ($70$ and $140$). In turn, the annealing-classical hybrid method offered by D-Wave Systems Inc. achieved a Q-score of $12{,}500$, which hints at gains offered by quantum annealing. The work of [18] also briefly discusses gate-based devices and their expected lower Q-score, mainly due to their low number of available qubits, and does not discuss photonic devices.
III The Q-score Framework and the Q-score Max-Clique
As mentioned in the previous section, the original Q-score definition leaves a couple degrees of freedom, the most notable ones being the allowed runtime and the amount of optimisation. Earlier work imposed specific values for these degrees of freedom [18], which yielded a completely set metric. In addition, the original Q-score definition makes specific choices for parameters which are inherently degrees of freedom, the most notable ones being the optimisation problem, the considered data set and the $\beta^{*}$-value, which measures the required accuracy. However, these are actually degrees of freedom which should be imposed by the use case. As (application-level) quantum metrics should optimally measure the performance of devices on applications, these parameters should also remain a degree of freedom in the definition of the metric.
This results in the definition of the Q-score framework, in which optimisation problem, data set, $\beta^{*}$-value, allowed runtime and allowed optimisation are all degrees of freedom which have to be chosen by the end user. Given a certain use case, the Q-score for this specific use case can then be defined by taking the following steps:
  1. 1.
    Choose (optimisation) problem that is most relevant to the use case. This problem needs to satisfy three requirements: firstly, it needs to be solvable for increasing problem sizes. Secondly, it needs to have a single value as an output, with a higher value corresponding to a better performance. Lastly, and most importantly, it needs to have a scalable approximation for the average value of a random and the optimal solution.
  2. 2.
    Choose a data set relevant to the use case on which devices can try to solve the considered problem.
  3. 3.
    Choose the required accuracy for the use case, resulting in a certain $\beta^{*}$-value.
  4. 4.
    Choose the allowed runtime for solving a single problem instance.
  5. 5.
    Choose the amount of optimisation that is allowed for solving a single problem instance. This amount should directly depend on the amount of (economic) resources that are available for the end user.
The resulting Q-score results in a metric that optimally measures the performance of a device at solving the use case.
It is good to emphasise that the Q-score framework in itself is nothing more than a way of combining the different degrees of freedom to yield a single number output. In the general definition, the Q-score framework depicts the largest problem size for which a device "on average" "significantly" outperforms a random approach at a certain problem on a certain data set, within the given time limit and allowed optimisation. In this definition, on average means averaged over 100 instances within the given data set. It should be noted that this number could also be seen as a degree of freedom, but it is not expected that changing this value will result in different metrics. A device significantly outperforms a random solution, if the cost $C(N)$ for a given problem size $N$ satisfies
$\beta(N):=\frac{C(N)-C_{rand}}{C_{max}-C_{rand}}>\beta^{*},$
where $C_{rand}$ and $C_{max}$ are the costs of a random and optimal solution of the considered problem.
To avoid confusion between different instantiations of the Q-score metric, we suggest to use the standalone name of Q-score only for the framework itself. To talk about a specific instantiation of the Q-score, the name Q-score accompanied with the optimisation problem should be used. If the other degrees of freedom are of relevance, these should be mentioned as well. Thus, the Q-score definition from [18] should be called the Q-score Max-Cut within this framework (on a data set of Erdös-Renyí ($N,\frac{1}{2})$-graphs with $\beta^{*}0.2$, 60 seconds runtime limit and no allowed (extra) optimisation).
III-A Q-score Max-Clique
To showcase the strength of the framework, a second instantiation of the Q-score framework will be given in this section, called the Q-score Max-Clique. The idea of this extension is to have a metric which can be be natively implemented on photonic quantum devices, in addition to gate-based quantum devices, quantum annealers and classical devices. This allows a fair comparison between all of these paradigms. The next section will show that this metric can indeed be used to benchmark all four of these paradigms.
@@ III-A Q-score Max-Clique
The Max-Clique problem is an established NP-hard problem in graph theory which aims to find the largest complete subgraph of a random graph of size $N$. A complete subgraph, or clique, is a subgraph in which every two vertices are connected by an edge. Just like the Max-Cut problem, the Max-Clique problem has a QUBO-formulation and the problem is hence equivalent to many optimisation problems, in particular to the Max-Cut problem [28]. The Q-score Max-Clique hence also serves as a metric on how well a device performs at general optimisation problems.
To define a Q-score using the Max-Clique problem, the Max-Clique problem needs to satisfy the three requirements as mentioned above, of which it clearly satisfies the first two. For the third requirement, a scalable approximation of the average random and optimal cost for solving the Max-Clique problem is required.
@@
The random cost $C_{rand}$ is defined as as the average clique size found by the following random naive approach: start with an empty sub-graph $G^{\prime}$; Take a random node $v\in G\setminus G^{\prime}$ and add this to $G^{\prime}$ if, $G^{\prime}\cup\{v\}$ is still a clique. If not, return the number of nodes in $G^{\prime}$ as the maximum clique size found by this random algorithm. For general Erdös-Rényi graphs $G(N,p)$, the probability that this naive algorithm finds a clique of size at least $i$ is
@@
For $p=\frac{1}{2}$, this converges quickly to the approximate value given by $C_{rand}=\mathbb{E}[X]=1.6416325$.
For the average cost $C_{max}$ of the optimal solution, there exist scalable expressions in literature [29], which say that this value scales as
@@
Note that this is an asymptotic result for $N\to\infty$ and might not be a good approximation for smaller $N$. With these two definitions, it can be concluded that the Max-Clique problem can indeed be used to define a Q-score metric.
In this work, we consider a specific instantiation in which the other degrees of freedom are also set. Firstly, as a data set, the same type of graphs as the original Q-score: $(N,\frac{1}{2})$-Erdös-Rényi graphs, and and the cost $C(N)$ is computed as the average clique size found over $100$ random instances of these graphs. Secondly, a value of $\beta^{*}=0.2$ is taken again, following [14]. Thirdly, the limit on the computation time is set to $60$ seconds per instance, as in [18]. If a device does not find an answer within the allocated 60 seconds, a clique size of $C_{rand}$ is assigned, corresponding to a value of $\beta(N)=0$. Finally, all algorithms are considered in an out-of-the-box fashion to restrict any additional optimisation steps.
All in all, the Q-score Max-Clique with these degrees of freedom is thus the largest problem size $N$ for which an unoptimised device obtains a cost $C(N)$ within 60 seconds for which
@@
IV Results
In this section the defined Q-score Max-Clique will be utilised to benchmark the different computational paradigms. It should be noted that the goal of this is not the benchmark itself, but rather to show that the metric can indeed be used to benchmark all the different platforms in a native way. First, a brief overview is given of how the Max-Clique problem can be solved on each of the quantum computing paradigms, as well as the devices for which the Q-score Max-Clique is computed.
@@ IV Results
  • Gate-based quantum computing: The same QUBO formulation can be used to solve Max-Clique instances using the Quantum Approximate Optimisation Algorithm (QAOA)[30]. This is a hybrid quantum-classical algorithm in which a certain cost function, depending on measurements of quantum states, is optimised with a classical optimiser. For each optimisation step, a new quantum circuit needs to be run. Currently, quantum devices are accessed through the cloud, and resources are shared with other users. Because of this, each set of quantum circuits needs to be sent as a job through the entire stack and queue, which is rather inefficient, especially in variational settings, such as QAOA. Solutions are being developed that provide users the ability to avoid latency by executing entire programs in containerised services, such as Qiskit Runtime [31]. Still, significant queue times remain and once scheduled, a single problem instance can still take around an hour to execute in practice. The actual computation time QAOA would take with direct access to dedicated hardware is hard to estimate. A methodology to reason about the execution-time of QAOA is developed in [32]. According to their estimates, candidate solutions for Max-Cut graphs with $500$ nodes can be generated in under three minutes, which is still above the $60$ seconds threshold.
    An extensive review of QAOA, including its various variants and the challenges related to their applicability, is provided in [33].
@@
In addition, the Q-score Max-Clique is also calculated for two classical solvers, namely qbsolv and simulated annealing, both offered by D-Wave Systems Inc. The qbsolv method solves smaller pieces of a QUBO with a tabu search algorithm [43, 44] and then combines those solutions to a full solution. The simulating annealing method is able to solve a QUBO by searching the solution space and slowly decreasing the probability of accepting worse solutions than the current optimum. Lastly, the Q-score Max-Clique is computed for the hybrid solver provided by D-Wave Systems Inc. This is an algorithm which combines classical and annealing approaches to solve QUBOs. The precise workings of this solver are however unknown, so the solver should be seen as a black-box device, which this may yield a somewhat unfair comparison. However, this is not an issue for this work, as the goal is to show that the Q-score Max-Clique can be used to benchmark all the different platforms, not to perform the comparison itself. This also demonstrates that the Q-score metric can be used to compare with various classical (approximate) algorithms.
Table I gives the Q-score of all considered backends.
Table I: Q-scores Max-Clique with a $60$ seconds time limit. The starred entries were obtained using more than 60 seconds.
@@
D-Wave Tabu search 4,900
D-Wave Simulated annealing 9,100